OpenStudy (anonymous):

Find x if y=-2 4x-6/y=15

OpenStudy (jtvatsim):

Do you have any ideas? What might you think to do first?

OpenStudy (anonymous):

I think subtract 4x in both sides

OpenStudy (jtvatsim):

So, if we do that, what will we have then?

OpenStudy (anonymous):

idk that's the part I don't get

OpenStudy (jtvatsim):

OK, good. We'll start there. You have 4x - 6/y = 15. By subtracting 4x from both sides we are saying: 4x - 6/y - 4x = 15 - 4x. The left side has a nice cancellation, 4x - 4x - 6/y = 15 - 4x 0 - 6/y = 15 - 4x -6/y = 15 - 4x This is what is left after subtracting 4x from both sides. Does that make sense so far?

OpenStudy (anonymous):

whats the answer though I need to find x

OpenStudy (jtvatsim):

Good! That's what I was going to ask next. The question wants us to find x. But we have a weird y in the equation. Do we know anything about y?

OpenStudy (anonymous):

you need to take out y

OpenStudy (jtvatsim):

That would be nice. Look at the question too. It seems to tell us that y = -2.

OpenStudy (jtvatsim):

So, we are allowed to replace any y's we see with a -2 instead.

OpenStudy (jtvatsim):

This means that -6/y = 15 - 4x is really the same as -6/-2 = 15 - 4x which is a little easier.

OpenStudy (jtvatsim):

Any questions on that part so far?